OpenStudy (anonymous):

solve for x x/6+x/12=3

OpenStudy (saifoo.khan):

Take LCD!!

hero (hero):

Hint: Multiply the first fraction by 2/2 to get: \(\large\frac{2x}{12} + \frac{x}{12} = 3\)

hero (hero):

Now combine fractions on the left side: \(\frac{3x}{12} = 3\) Then multiply both sides by 12 to get \(3x = 36\) Finish it

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

thanks hero